OpenStudy (anonymous):

How to find the Domain of f(x) = tan(x) - 3 ?

OpenStudy (anonymous):

\[domain={R-n \frac{ \pi }{ 2 }} ,where n is an integer.\]

OpenStudy (anonymous):

Does the R "All Real Numbers"?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you think you could explain how to get the answer? I will be needing to do this for a test. I'm comfortable finding domain and range for sine and cosine.

OpenStudy (anonymous):

\[\tan x is notdefined at atx=.....\frac{- \pi }{2},\frac{ \pi }{ 2 },\frac{ 3\pi }{ 2 },.....\] \[so domain is all real numbers except at x=.....,\frac{ \pi }{ 2 },\frac{ \pi }{ 2 },\frac{ 3 \pi }{2 },....\]